The sum of first 4 terms of an A.P is 26 and sum of first 8 terms is 100. Find the first 28 terms

Grade:11

1 Answers

Arun
5 years ago
Dear Bipin
 
S4 = 26
Hence
(4/2)( 2a + 3d) = 26
2a + 3d = 13
S8 = 100
(8/2)(2a + 7d) = 100
2a + 7d = 25
Solve both the equations
4d = 12
d = 3
Hence
a = 2